# A Video is Worth 10,000 Words: Training and Benchmarking with Diverse Captions for Better Long Video Retrieval

### v1

@@ Abstract
Existing long video retrieval systems are trained and tested in the paragraph-to-video retrieval regime, where every long video is described by a single long paragraph. This neglects the richness and variety of possible valid descriptions of a video, which could be described in moment-by-moment detail, or in a single phrase summary, or anything in between. To provide a more thorough evaluation of the capabilities of long video retrieval systems, we propose a pipeline that leverages state-of-the-art large language models to carefully generate a diverse set of synthetic captions for long videos. We validate this pipeline’s fidelity via rigorous human inspection. We then benchmark a representative set of video language models on these synthetic captions using a few long video datasets, showing that they struggle with the transformed data, especially the shortest captions. We also propose a lightweight fine-tuning method, where we use a contrastive loss to learn a hierarchical embedding loss based on the differing levels of information among the various captions. Our method improves performance both on the downstream paragraph-to-video retrieval task (+1.1% R@1 on ActivityNet), as well as for the various long video retrieval metrics we compute using our synthetic data (+3.6% R@1 for short descriptions on ActivityNet).
**footnotetext: Work performed during internship with SRI International.
@@ 1 Introduction
If a picture is worth 1,000 words, then a video is worth 10,000. Consider the variety of possible captions for a video, illustrated in Figure 1, which describe the video at varying levels of detail. These can vary substantially in terms of semantics and structure. Nevertheless, a video retrieval system ought to be able to map any of these captions to the video, whether they are short, long, abstract, or even describe only part of the video. We thus seek to redefine the video retrieval problem, where a system, given some text, must retrieve a video.
@@
Video retrieval has long been a important interdisciplinary task that connects the state-of-the-art in natural language processing and computer vision [25, 34]. At its core, solving the problem requires a system that understands both video and text. However, current video retrieval literature often considers just short clips, which cannot be described by such a variety of captions, and thus obscures the problem. Increasingly more work does focus on long videos, but it uses only full paragraphs for retrieval [7, 6], neglecting to consider the rich space of valid video captions. However, the real world includes long videos with multiple events, and these admit many vague, abstract, or partial descriptions a person (e.g., doing video search) might give. In practice, the strict mapping of a single caption to a single video in instance-based video retrieval neglects both the potential ambiguities of the caption as well as the fact it is valid to describe the same video in a variety of ways [48].
To address this we formulate the 10k Words problem, a novel video retrieval setting which includes diverse descriptions generated for long videos with multiple events. This problem identifies key axes of variation, including simplification, summarization, and duration, that we use to curate pools of captions with non-trivial differences in structure and semantics. The problem introduces challenging ambiguities, since some captions will not mention all the details that distinguish a video from similar, related videos.
We instantiate the 10k Words problem by building 3 datasets, ActivityNet10k, QuerYD10k, and LF-VILA10k, by augmenting existing datasets [17, 37, 45] with diverse captions. These augmentations are only possible given the flexibility and accuracy of recent large language models (LLMs) [8], which we combine with some simple automatic manipulations to synthesize the diverse 10k Words datasets. We provide a thorough definition of the 10k Words problem and our data generation pipeline in Section 2.
These datasets can help us detect failures of existing models to capture the space of text descriptions as well as help us to mitigate those failures, and we show both. For detection, we consider a representative set of state-of-the-art video models and show that they struggle to adequately solve the 10k Words problem in Section 3, struggling especially with short, summary-style captions. We then demonstrate the effectiveness of a simple mitigation strategy that uses 10k datasets to augment standard datasets during training. This can provide an inexpensive boost to performance on both the 10k datasets and the original standard datasets, or be used to increase data efficiency. However, simple data augmentation does not fully leverage the knowledge that multiple possible captions correspond to the same video despite meaningful structural and semantic differences.
We thus propose an approach that uses 10k captions to guide the text embedding space. Prior work has pointed out that some paragraph captions may already be valid descriptions of multiple videos [48], and this is only more frequent with our synthetic short summary captions. To help our model address this we finetune it to not only bring synthetic caption features and real video features close to each other, but also to align a projection of the synthetic text features with real text and real video features. The projection step allows the model to learn embeddings that account for the different levels of information between the synthetic caption, the real caption, and the real video. We achieve SOTA performance on 10k Words while also boosting performance on the standard paragraph-to-video retrieval task. In summary, we contribute the following:
  • We formulate and propose the 10k Words problem, a framework for characterizing the broad spectrum of valid descriptions for long videos.
  • We create ActivityNet10k, QuerYD10k, and LF-VILA10k by using our flexible data generation pipeline.
  • We benchmark SOTA models in a zero-shot fashion, and reveal that they struggle on the 10k Words problem.
  • We leverage 10k data and lightweight contrastive losses to achieve SOTA performance for 10k Words, with a +1.1% R@1 improvement on standard ActivityNet retrieval and +2.8% R@) improvement on ActivityNet10k.
2 10k Words Dataset
2.1 10k Definition and Generation
@@
Table 2: Automatic dataset statistics. We show the average change in unique nouns and verbs, as well as word count and length for the different dimensions of our 10k Words ActivityNet datasets.
@@
Metric
Source Short Medium Full Length Primary Secondary University S and P S and S S and U
$\Delta$ Nouns
-5.37 -1.69 0.06 -1.16 0.97 3.84 -6.20 -5.83 -5.41
$\Delta$ Verbs
-5.01 -1.97 -0.77 -0.65 0.90 1.95 -5.19 -5.02 -4.84
Word Count
49.77 8.77 29.29 37.39 43.54 48.31 56.13 8.50 9.28 10.75
Word Length
5.09 6.10 5.40 5.51 4.97 5.49 5.97 5.27 5.75 6.10
We now describe the procedure for creating a dataset of videos and natural language description queries such that the queries cover many of the possible ways of describing a video. We start with an existing dataset of annotated videos and enrich the set of descriptions in the dataset, so it covers more of the possible ways to describe the videos. To cover the broadest possible spectrum of natural language queries for a video we start by defining three axes along which a video’s description can vary: duration, summarization, and simplification. The first dimension, duration, refers to how many of the events in a video are described by a given query. The last two dimensions cover different ways of using language to describe the same video. Prior to LLMs these last two dimension would have been difficult to simulate effectively and reliably, and perhaps would have required expensive human annotations. However, we are able to effectively prompt LLM to gather such data [8, 57]. For examples of these dimensions, see Table 1.
More precisely, consider a long video with $E$ events ${e_{1},e_{2},\ldots,e_{E}}$. Each event $e_{i}$ has a corresponding short video clip $v_{i}$ and natural language description of that clip $t_{i}$ with the set of clips and texts being denoted $V$ and $T$. As such, the original long caption could be a paragraph, long sentence, or, more typically, the concatenation of video segment captions. This is typical for the datasets shown in Table 3. Each axis of the 10k problem implements a function that takes such a video with event segmentation and descriptions as input and outputs a new augmented version of the same video with a new set of segments and descriptions. We now cover each of these functions.
Summarization. Descriptions of videos can vary in length, where at one extreme they describe every detail in the video, at the other they briefly describe the main idea leaving out some significant details, and in between relevant details are progressively grouped and redundant elements are pruned. At one end of this spectrum a video retrieval model must be able to parse details and at the other end it must be able to understand a gestalt. To augment a video on this axis we prompt an LLM with the ground truth descriptions $T$ (concatenated) and instruct it to generate summaries. If the concatenated description has $L$ words then we ask the LLM to generate three summaries with $\lfloor{L}\cdot\frac{l}{7}\rfloor$ words each for $l\in\{1,4,7\}$. At full length ($l=7$) this should just re-phrase the concatenated caption, but at smaller lengths the LLM must leave out information. We observe that GPT-3.5***gpt-3.5-turbo-0613 is able to achieve close to the desired word count most of the time. This only changes $T$, leaving $E$ and $V$ unchanged. We share exact prompts in the appendix.
Table 3: Base Datasets which we use for our synthetic caption generation pipeline. We use only the ‘train’ and ‘val-1’ splits of ActivityNet, and do some additional filtering for extremely long outlier captions from LF-VILA and QuerYD.
Dataset # Videos Video Len (s) Text Len (w) Duration (h)
ActivityNet [27] 14926 117.5 49.8 487
LF-VILA-7k [54, 45] 7020 203.9 155.4 397
QuerYD [37] 2474 264.3 203.8 181
Simplification. Descriptions of videos can vary in terms of their conceptual simplicity, where an idea could be described at the level of a college graduate, or else simplified for a kindergartener, and a good retrieval model should map all these descriptions to the same video. We capture this dimension by providing an LLM with the same ground truth description as for summarization and instruct it to output a simplified version. This is done for three levels of reading comprehension described to the LLM as “primary school”, “secondary school”, or “university” reading level. This augmentation only modifies $T$, leaving $E$ and $V$ unchanged. We share exact prompts in the appendix.
Duration. Descriptions of videos can be partial, intending to cover only a segment of the video, but the video should still be retrieved when these are used as queries. In our dataset we implement this by choosing a contiguous subset of events $\tilde{E}={e_{i},\ldots,e_{j}}$ with start and end index $i$ and $j$. The corresponding set of video clips $\tilde{V}$ and captions $\tilde{T}$ are selected to create the augmented video.
@@
10k Datasets. We combine these axes to construct 10k versions of ActivityNet, QuerYD, and LF-VILA-7k (Table 3). We construct our 10k Words datasets by taking the per-segment captions available for the datasets described in Table 3 and feed them to GPT-3.5 with relevant prompts. Starting from each video in a base dataset like ActivityNet, we include 11 captions for each video: 1 full caption (original ground truth paragraph), 3 captions for the levels of simplification (elementary, intermediate, and university), 3 captions for the levels of summarization (short, medium, and long), 3 captions that combine summarization and simplification by generating simplifications for the short summaries, and 1 caption corresponding to a random subset of the original video segments by duration augmentation. We show examples for summarization, simplification, and partial captions, and introduce relevant shorthand, in Table 1.
We refer to the 10k Words version of LF-VILA-7k, a sample from the original LF-VILA [45], as LF-VILA10k. We use all of QuerYD as a validation set, since the initial small size of its validation set makes it challenging to distill useful insights, and create QuerYD10k. We create ActivityNet10k for ActivityNet. We provide details on LLM prompts and costs in the appendix.
2.2 Dataset Analysis
@@
We provide some fine-grained statistical measures to examine the nature of our generated data in Section 2.2.1. We also perform a study on a sample of our data using human annotators to further validate the claims regarding our data and ensure that it is free from undesirable artifacts (e.g. hallucinations) in Section 2.2.2.
2.2.1 Automatic Analysis
@@
First, we perform an automatic analysis of our generated data. For the sake of brevity we focus on ActivityNet, with metrics for other datasets provided in the supplement. From Table 2, note that summarization and elementary level simplification tend to remove nouns and verbs ***Extracted using https://spacy.io part-of-speech tagging, while intermediate and university levels tend to add nouns and verbs. Also note the word counts, where summarization, as expected, reduces the average number of words, while simplification to elementary level reduces the average length of words.
@@
Table 8: Text to Video retrieval performance for our benchmark on ActivityNet. First we reproduce results for standard paragraph to video retrieval. Then, we give the average performance on short 10k Words captions, long 10k Words captions, and partial captions. We use the standard recall at top-1 metric (R@1) as well as the average of recall at top-1/5/10 (Avg. Recall). We explain how we aggregate 10k Words captions as Full, Short, Long, and Partial in Section 3.1. We explain the finetuning methods in Section 4.
@@
ANet Full QuerYD Full LF-VILA Full ANet10k All ANet10k Short ANet10k Long ANet10k
@@
finetune Domain 58.8 77.7 - - 95.9 98.6 42.0 62.8 29.9 51.4 53.8 73.8 43.2 64.5
10k words 59.3 78.1 - - 97.4 99.1 44.3 65.2 32.3 54.8 56.2 75.5 44.3 65.9
Ours 59.9 78.5 - - 97.5 99.1 44.8 66.1 33.5 56.3 56.2 75.9 44.7 66.2
2.2.2 Annotator Analysis
To validate the fidelity and utility of our captions we recruit 15 human annotators to examine our captions in an IRB approved study. We design a survey that consists of 3 sections, according to the properties of the data we wish to examine. In the first section, we analyze whether the LLM-generated captions preserve original meaning. In the second section, we ensure that when the LLM performs the simplification in a manner that is meaningful to humans. In the third section, we verify the extent to which hallucinations occur in the LLM-generated captions. Each section has 5 questions. We divide our annotators into groups of 3, to allow for analysis of inter-annotator agreement, and thus distribute 5 versions of the survey, covering a sample of 75 videos from the validation set of ActivityNet. Next, we provide more detail regarding the design of the survey and results for each section. As evidence of the survey’s validity, we show inter-annotator agreement in Table 7. For full details, including exact surveys, please see the appendix.
Meaning Preservation. For each question in this section, we randomly sample one real caption, and assign it to be the “ground truth” caption. We then sample 3 generated captions – one generated from the “ground truth” caption, one generated from the “ground truth” caption’s nearest neighbor caption***Maximum cosine similarity between “ground truth” captions using OpenAI text-embedding-ada-002 embedding., and one generated from a random unrelated caption. We ask the annotator to determine, for each of the 3 generated captions, whether they believe it describes the same video as the ground truth caption. We observe, in Table 6, that the generated caption for any given source caption is very consistently judged to be from the same video, with neighbors being determined to come from the same video to a lesser extent.
Simplification Validation. For each question in this section, we randomly sample one real caption and show the annotator the “primary school”, “secondary school”, and “university” captions generated by the LLM. Then we ask the annotator to rank them from most to least complex. In Table 6, we find very little ambiguity in the simplification rankings. Annotators consistently judge “university” to be the most complex, and “primary school” to be the least complex.
Hallucination Prevalence. For each question in this section, we sample some real caption and one of its generated captions, either the full length summary, or one of the full length simplification captions. We then use spaCy***https://spacy.io part-of-speech tagging to extract the nouns and verbs which appear in the generated caption but not the original. We then present both captions to the annotator, and for up to 3 of the potentially hallucinated words, we ask whether or not they change the meaning of the original caption. We find, in Table 6, that of the new words for the generated captions, annotators tend to judge that they typically correspond to entities and actions that are already depicted in the source captions. This, along with the results from the first section of the study, suggest that the prevalence and impact of potential hallucination is quite limited.
All this automatic and human analysis leads us to conclude that the captions we generate in our novel paradigm are diverse, robust, and, as we will continue to demonstrate, very useful.
3 Benchmark and Analysis
Figure 2: We compute the difference in R@1 between retrieval with full paragraph captions, compared to short (s), medium (m), long(l), elementary (e), intermediate(i), university (u), and partial captions and show the difference as a percentage relative to the full paragraph R@1. Each bar is the average of these percentages for all 3 datasets. Notice that while deltas are consistent between medium and long captions, there is a drop off for short captions and partial captions, demonstrating that a key facet of this problems difficulty is the introduction of ambiguity due to missing information.
Figure 3: We measure the retrieval for the datasets across all zero-shot models and captions. We plot the number of videos which are retrieved as the top-1 result for the given number of model-caption pairs. At $x=0$, this is the number of videos which are not retrieved by many model, for any caption. At $x=40$, this is how many videos are retrieved as the first result for all 4 models, for all 10 possible captions.
3.1 Preliminaries
Models
@@ 3.1 Preliminaries
For our 10k Words benchmark, we evaluate the performance of VideoCLIP [50], Frozen [7], InternVideo [46], and COSA [12] as a set of representative video-language models. For COSA we use the ‘itm’ retrieval and for InternVideo we use the dual softmax. We evaluate these in a zero-shot manner to show 10k data is difficult for existing models, and show that both finetuning and our proposed approach can improve performance in Section 4.
Experiment Details
@@
We run our experiments on nodes containing between 1 and 4 GTX 2080Ti, RTX A5000, and RTX A6000 GPUs, depending on the demands of each model. When training, we follow the settings provided in the publicly available code of the models we chose. Please see our code (with our dataloaders and other relevant modifications) for reproducibility.
@@
Metrics
We primarily use text-to-video recall @ K (R@K) to measure performance. Given a list of text queries and video targets relative to a database of videos to be retrieved, R@K measures the percentage of queries for which the ground truth target was retrieved at rank K. We use this to measure overall performance, but also to measure subset performance on Summarization and Simplification by considering the appropriate subsets of captions.
To measure performance on the Duration axis we consider whether the partial and full captions can retrieve the full length video. The “Full” setting measures how often the full caption (f) retrieves the video at rank K or better, which represents performance as measured by the standard datasets. Since we use the standard ActivityNet settings, these can be compared with numbers from other papers; however, since we use unique splits for QuerYD and LF-VILA, these numbers are not comparable. The “Partial” setting measures how often the partial caption (p) retrieves the same full length video. The “Short” setting measures performance of full length video retrieved by short captions including short summarization (s) and simplifications of it (s+e, s+i, s+u). Similarly, we also report performance on the “Long” setting which include long summarization (l) and simplifications of it (l+e, l+i, l+u). The “All” setting is an average of the Partial, Short, and Long, weighted corresponding to the number of captions that comprise each.
3.2 Zero-Shot Results
@@
We provide the zero-shot benchmarking results in Table 8, with remaining results on the LF-VILA and QuerYD in Table 9. The methods that perform best for Full paragraphs also tend to perform best for the Long, Short, and Partial captions. Notably, there is only a minor gap in retrieval performance between the Full and Long captions, with a larger difference between Full and Partial captions, and a significant drop for the Short captions. We also show using Figure 2 that the shortest summaries are the hardest to retrieve. Note that the differences in Figure 2 are computed as a percentage of the Full R@1. With this in mind, we see that COSA seems to be by far the least robust to 10k Words data, with the largest relative changes in performance. By contrast, VideoCLIP and especially Frozen are often benefited by the 10k Words data, particularly when the axis is rewording the caption (e, i, u) rather than removing it (s, p).
With these interesting performance discrepancies in mind, we leverage our suite of models and 10k Words data as a way to evaluate the retrieval difficulty of each dataset. To do so, we consider all 10 possible captions (we exclude Partial) and a model as 10 model-caption pairs, so with 4 models, we have 40 pairs. We then measure, for each video in a dataset, how often (for the 40 pairs) it is retrieved as the first-ranked result. We plot this in Figure 3 and find that LF-VILA tends to be the easiest dataset, with QuerYD having a significant portion of videos that are not retrieved correctly no matter which model and caption are used.
4 Long Video Finetuning
Figure 4: We perform contrastive finetuning for retrieval with video-caption pairs.  One key novelty is the project we propose for 2 projection losses, between projected 10k  text features and both ground truth text features as well as video features, respectively.  The other key novelty is the data augmentation, where we compute standard contrastive loss but each caption is sampled  randomly from the complexity distribution for that entity, according to a mixing ratio, $\eta$.
In this section we propose a method for finetuning pre-trained models to retrieve videos from any relevant caption. We then highlight results.
4.1 Method

### v2

@@ Abstract
Existing long video retrieval systems are trained and tested in the paragraph-to-video retrieval regime, where every long video is described by a single long paragraph. This neglects the richness and variety of possible valid descriptions of a video, which could range anywhere from moment-by-moment detail to a single phrase summary. To provide a more thorough evaluation of the capabilities of long video retrieval systems, we propose a pipeline that leverages state-of-the-art large language models to carefully generate a diverse set of synthetic captions for long videos. We validate this pipeline’s fidelity via rigorous human inspection. We use synthetic captions from this pipeline to perform a benchmark of a representative set of video language models using long video datasets, and show that the models struggle on shorter captions. We show that finetuning on this data can both mitigate these issues (+2.8% R@1 over SOTA on ActivityNet with diverse captions), and even improve performance on standard paragraph-to-video retrieval (+1.0% R@1 on ActivityNet). We also use synthetic data from our pipeline as query expansion in the zero-shot setting (+3.4% R@1 on ActivityNet). We derive insights by analyzing failure cases for retrieval with short captions.
@@ 1 Introduction
If a picture is worth 1,000 words, then a video is worth 10,000. Consider the variety of possible captions that can describe just one video (Figure 1). Although they can vary substantially in semantics and structure, video-language models ought to be able to match all of these captions with the video they describe.
In this paper we show that existing approaches fail to model the variety of captions and show how they can be improved in the context of video retrieval. At its core, video retrieval requires not just a system that understands video and text, but also how minor differences between videos in a dataset make them unique. However, video retrieval literature traditionally considers just short clips [25, 35], which cannot be described by such a variety of captions, and thus obscures the problem. Increasingly more works have focused on long videos with multiple events, but it uses only full paragraphs for retrieval [7, 6], neglecting the rich space of valid captions. While even existing captions can be ambiguous [48], they still do not include vague, abstract, or partial descriptions a user (e.g., doing video search) might give. This means current video retrieval datasets do not measure real world performance, where captions can be ambiguous, vary in semantics and style, and can describe long complex videos.
To address this we formulate the 10k Words benchmark, a novel video retrieval setting which includes diverse descriptions generated for long videos with multiple events. We identify key axes of variation, including simplification, summarization, and duration, then use them to curate pools of captions with non-trivial differences in structure and semantics. The benchmark introduces challenging ambiguities, since some captions will not mention all the details that distinguish a video from similar, related videos. We instantiate this benchmark by augmenting existing datasets [17, 37, 45] with diverse captions, creating ActivityNet10k, QuerYD10k, and LF-VILA10k (borrowing 10k from the idea that “a video is worth 10,000 words,” and we work towards that richness of description with our diverse sythetic captions). These augmentations are only possible given the flexibility and accuracy of recent large language models (LLMs) [8], which we combine with some simple automatic manipulations to synthesize the diverse 10k Words datasets as described in Section 2.
These proposed datasets can help us detect failures of existing models to capture the space of text descriptions as well as help us to mitigate those failures, and we show both. For detection, we consider a representative set of state-of-the-art video models and show that they struggle to adequately solve the 10k Words problem in Section 3, struggling especially with short, summary-style captions. We then demonstrate the effectiveness of a simple mitigation strategy that uses 10k datasets to augment standard datasets during training. This can provide an inexpensive boost to performance on both the 10k datasets and the original standard datasets, or be used to increase data efficiency. We also investigate inference time improvement, showing how query expansion can benefit a pre-trained model without finetuning. An LLM can be used during inference to improve retrieval performance by generating multiple queries for the same video. We achieve SOTA performance on 10k Words while also boosting performance on the standard paragraph-to-video retrieval task.
Finally, in Section 5 we analyze failure cases to understand whether shorter captions are truly ambiguous or not. We find some cases our not ambiguous, indicating our models have room for improvement on our 10k benchmark. In summary, we contribute the following:
  • We instantiate the 10k Words benchmark, a framework for characterizing the broad spectrum of valid descriptions for long videos, by creating ActivityNet10k, QuerYD10k, and LF-VILA10k with a flexible LLM-based pipeline.
  • We evaluate SOTA models in a zero-shot fashion, and reveal that they struggle on the 10k Words benchmark.
  • We leverage 10k data for an improvement of +3.4% R@1 on zero-shot standard ActivityNet retrieval (without finetuning) and +2.8% R@1 on ActivityNet10k retrieval (with finetuning), which is SOTA 10k Words performance.
  • We show that despite the ambiguity of shorter captions, SOTA models still fail in non-ambiguous cases.
2 10k Words Benchmark and Datasets
2.1 10k Definition and Generation
Given an existing dataset of videos and corresponding descriptions, we create a 10k version of the dataset by enriching the set of descriptions to cover more possible ways to describe the videos. Existing datasets like those in Table 2 often take a long video and annotate $E$ events ${e_{1},e_{2},\ldots,e_{E}}$ individually. Each event $e_{i}$ has a corresponding short video clip $v_{i}$ and is annotated with a natural language description of that clip $t_{i}$, with the set of clips and texts for a given video being denoted $V$ and $T$. As such, the original long caption could be a paragraph, long sentence, or, more typically, the concatenation of video segment captions, which is then treated as a paragraph. To cover the broadest possible spectrum of natural language queries for a video we start by defining three augmentation axes along which a video’s description can vary: duration, summarization, and simplification. Duration refers to how many of the events in a video are described by a given query, while summarization and simplification cover different ways of using language to describe the same video. For each axis we implement a function that takes a video with event segmentation and descriptions as input and outputs a new augmented version of the same video with a new set of segments and descriptions. Prior to LLMs, summarization and simplification would have been difficult to simulate effectively and reliably, and perhaps would have required expensive human annotations. However, now we are able to effectively prompt LLM to gather such data [57, 8]. Next we discuss the prompts we design for each augmentation axis, with example synthetic notations in Table 1.
@@
Table 2: Datasets which we create with our synthetic caption generation pipeline. We use only the ‘train’ and ‘val-1’ splits of ActivityNet Captions, and do some additional filtering for extremely long outlier captions from LF-VILA and QuerYD.
Dataset Source Dataset # Videos Video Len (s) Text Len (w)
ActivityNet10k ActivityNet Captions [27] 14926 117.5 49.8
LF-VILA10k LF-VILA [54, 45] 7020 203.9 155.4
QuerYD10k QuerYD [37] 2474 264.3 203.8
Summarization. Descriptions of videos can vary in length. While at one extreme they describe every detail in the video, at the other they briefly describe the main idea, leaving out some significant details. In between the two extremes, relevant details are progressively grouped and redundant elements are pruned. At one end of this spectrum a video retrieval model must be able to parse details and at the other end it must be able to understand a gestalt. To augment a video on this axis we prompt an LLM with the ground truth descriptions $T$ (concatenated) and instruct it to generate summaries. If the concatenated description has $L$ words, then we ask the LLM to generate three summaries with $\lfloor{L}\cdot\frac{l}{7}\rfloor$ words each for $l\in\{1,4,7\}$. At full length ($l=7$) this should just re-phrase the concatenated caption, but at smaller lengths the LLM must leave out information. We observe that GPT-3.5111gpt-3.5-turbo-0613 is able to achieve close to the desired word count most of the time. This only changes $T$, leaving $E$ and $V$ unchanged.
Simplification. Descriptions of videos can vary in terms of their conceptual simplicity, where an idea could be described at the level of a college graduate, or else simplified for a kindergartener, and a good retrieval model should map all these descriptions to the same video. We capture this dimension by providing an LLM with the same ground truth description as for summarization and instruct it to output a simplified version. This is done for three levels of reading comprehension described to the LLM as “elementary”, “intermediate”, or “university” reading level. This only modifies $T$, leaving $E$ and $V$ unchanged.
Duration. Descriptions of videos can be partial, intending to cover only a segment of the video, but the video should still usually be retrieved when these are used as queries (see more about ambiguity in Section 5). In our dataset we implement this by choosing a contiguous subset of events $\tilde{E}={e_{i},\ldots,e_{j}}$ with start and end index $i$ and $j$. The corresponding set of video clips $\tilde{V}$ and captions $\tilde{T}$ are selected to create the augmented video.
10k Datasets. We combine these axes to construct 10k versions of ActivityNet Captions (ActivityNet), QuerYD, and LF-VILA (Table 2). We construct our 10k Words datasets by taking the per-segment captions available for the datasets described in Table 2 and feed them to GPT-3.5 with relevant prompts. Starting from each video in a base dataset like ActivityNet, we include 11 captions for each video: 1 full caption (original ground truth paragraph), 3 captions for the levels of simplification (elementary, intermediate, and university), 3 captions for the levels of summarization (short, medium, and long), 3 captions that combine summarization and simplification by generating simplifications for the short summaries, and 1 caption corresponding to a random subset of the original video segments by duration augmentation. We show examples and introduce relevant shorthand in Table 1.
We refer to the 10k Words version of LF-VILA, a sample from the original LF-VILA [45], as LF-VILA10k. We use all of QuerYD as a validation set, since the initial small size of its validation set makes it challenging to distill useful insights, and create QuerYD10k. We create ActivityNet10k for ActivityNet. We provide details on LLM prompts and costs in the appendix.
2.2 Dataset Analysis
@@
Table 3: The average change in unique nouns and verbs relative to the ground truth, as well as word count and length for the different dimensions of ActivityNet10k, vs. ActivityNet Captions [27].
@@
Metric Source Short Medium Full Length Elementary Intermediate University S and E S and I S and U
$\Delta$ Nouns -5.37 -1.69 0.06 -1.16 0.97 3.84 -6.20 -5.83 -5.41
$\Delta$ Verbs -5.01 -1.97 -0.77 -0.65 0.90 1.95 -5.19 -5.02 -4.84
Word Count 49.77 8.77 29.29 37.39 43.54 48.31 56.13 8.50 9.28 10.75
Word Length 5.09 6.10 5.40 5.51 4.97 5.49 5.97 5.27 5.75 6.10
@@
We provide some fine-grained statistical measures to examine the nature of our generated data (Automatic Analysis). We also perform a study on a sample of our data using human annotators to further validate the claims regarding our data and ensure that it is free from undesirable artifacts (Annotator Analysis).
@@
Automatic Analysis. For the sake of brevity we focus on ActivityNet, with metrics for other datasets provided in the appendix. From Table 3, note that summarization and elementary level simplification tend to remove nouns and verbs  [^2] Extracted using https://spacy.io part-of-speech tagging, while higher reading levels tend to add nouns and verbs. Also note the word counts, where summarization, as expected, reduces the average number of words, while simplification to elementary level reduces average word length.
Annotator Analysis. To validate the fidelity and utility of our captions we recruit 15 human annotators to examine our captions in an IRB approved study. We design a survey that consists of 3 sections, according to the properties of the data we wish to examine. In the first section, we analyze whether the LLM-generated captions preserve original meaning. In the second section, we ensure that when the LLM performs the simplification in a manner that is meaningful to humans. In the third section, we verify the extent to which hallucinations occur in the LLM-generated captions. Each section has 5 questions. We divide our annotators into groups of 3, to allow for analysis of inter-annotator agreement, and thus distribute 5 versions of the survey, covering a sample of 75 videos from the validation set of ActivityNet. Next, we provide more detail regarding the design of the survey and results for each section. As evidence of the survey’s validity, we show inter-annotator agreement in Table 7. For full details, please see the appendix.
Meaning Preservation. For each question in this section, we randomly sample one real caption, and assign it to be the “ground truth” caption. We then sample 3 generated captions – one generated from the “ground truth” caption, one generated from the “ground truth” caption’s nearest neighbor caption  [^3] From cosine similarity between captions using OpenAI’s text-embedding-ada-002., and one generated from a random unrelated caption. We ask the annotator to determine, for each of the 3 generated captions, whether they believe it describes the same video as the ground truth caption. Table 6 shows that the synthetic caption is very consistently judged to be from the same video as its source caption, unlike neighbor and random captions.
Simplification Validation. For each question in this section, we randomly sample one real caption and show the annotator the “elementary”, “intermediate”, and “university” captions generated by the LLM. Then we ask the annotator to rank them from most to least complex. In Table 6, we find very little ambiguity in the simplification rankings. Annotators consistently judge “university” to be the most complex, and “elementary” to be the least complex.
Hallucination Prevalence. For each question in this section, we sample some real caption and one of its generated captions, either the full length summary, or one of the full length simplification captions. We then use spaCy part-of-speech tagging to extract the nouns and verbs which appear in the generated caption but not the original. We then present both captions to the annotator, and for up to 3 of the potentially hallucinated words, we ask whether or not they change the meaning of the original caption. We find, in Table 6, that of the new words for the generated captions, annotators tend to judge that they typically correspond to entities and actions that are already depicted in the source captions. This, along with the results from the first section of the study, suggest that the prevalence and impact of potential hallucination is quite limited.
3 Benchmark Results
@@
Table 8: Text to Video retrieval performance for our benchmark on ActivityNet. First we reproduce results for standard paragraph to video retrieval. Then, we give the average performance on short 10k Words captions, long 10k Words captions, and partial captions. We use the standard recall at top-1 metric (R@1) as well as the average of recall at top-1/5/10 (Avg. Recall). We explain how we aggregate 10k Words captions as Full, Short, Long, and Partial in Section 3. We explain the finetuning in Section 4.
@@
ANet Full QuerYD Full LF-VILA Full ANet10k All ANet10k Short ANet10k Long ANet10k Partial
@@
finetune Domain 59.1 77.7 - - 95.9 98.6 43.0 64.2 29.1 51.5 54.2 74.5 43.2 64.5
Ours 60.1 78.7 - - 97.4 99.1 45.8 67.6 32.9 57.1 56.5 76.5 44.3 65.9
From Section 2.2, we conclude that the captions we generate in our paradigm are diverse and robust. In this section we demonstrate that they are useful for benchmarking the text-to-video retrieval performance of video-language models.
@@
Models For our 10k Words benchmark, we evaluate the performance of VideoCLIP [50], Frozen [7], InternVideo [46], and COSA [12] as a set of representative video-language models. For COSA we use the ‘itm’ retrieval and for InternVideo we use the dual softmax.
@@
Experiment Details We run our experiments on nodes containing between 1 and 4 GTX 2080Ti, RTX A5000, and RTX A6000 GPUs, depending on the demands of each model. When training, we follow the settings provided in the publicly available code of the models we chose.
Metrics We use text-to-video recall @ K (R@K) to measure performance. Given a list of text queries and video targets relative to a database of videos to be retrieved, R@K measures the percentage of queries for which the ground truth target was retrieved at rank K. Avg. R averages R@1, R@5, and R@10.
To measure performance on the Duration axis we consider whether the partial and full captions can retrieve the full length video. The “Full” setting measures how often the full caption (f) retrieves the video at rank K or better, which represents performance as measured by the standard datasets. Since we use the standard ActivityNet settings, these can be compared with numbers from other papers; however, since we use unique splits for QuerYD and LF-VILA, these numbers are not comparable. The “Partial” setting measures how often the partial caption (p) retrieves the same full length video. The “Short” setting measures performance of full length video retrieved by short captions including short summarization (s) and simplifications of it (s+e, s+i, s+u). Similarly, we also report performance on the “Long” setting which include long summarization (l) and simplifications of it (l+e, l+i, l+u). The “All” setting is an average of Partial, Short, and Long, weighted by the number of caption types for each.
@@
We provide the zero-shot benchmarking results in Table 8, with remaining results on the LF-VILA and QuerYD datasets in Table 9. The methods that perform best for Full paragraphs also tend to perform best for the Long, Short, and Partial captions. Notably, there is only a minor gap in retrieval performance between the Full and Long captions, with a larger difference between Full and Partial captions, and a significant drop for the Short captions. We also see that COSA seems to be by far the least robust to 10k Words data, with the largest relative changes in performance. By contrast, VideoCLIP and especially Frozen are often benefited by the 10k Words data, particularly when the axis is rewording the caption (e, i, u) rather than removing information from it (s, p).
4 Improving Performance
In this section we present baseline results where we finetune pre-trained models to retrieve videos from text. We then explore two main ways to leverage our data to improve these results. The first is at training time – with no extra cost in terms of parameters, iterations, or FLOPS, we can train with synthetic captions to improve retrieval results. The second is at inference time – we can leverage the 10k prompts as a form of query expansion, and aggregate the retrievals across equivalent 10k captions.
4.1 Training-time Improvements